# Enhancing Clinical Trial Patient Matching through Knowledge Augmentation and Reasoning with Multi-Agents

### v2

@@
Abstract.
Matching patients effectively and efficiently for clinical trials is a significant challenge due to the complexity and variability of patient profiles and trial criteria. This paper introduces Multi-Agents for Knowledge Augmentation and Reasoning (MAKAR), a novel framework that enhances patient-trial matching by integrating domain-specific knowledge with structured reasoning. MAKAR consists of two key modules: the Augmentation Module and the Reasoning Module. The Augmentation Module enriches trial criteria by incorporating detailed explanations of relevant concepts to ensure clarity and completeness. The Reasoning Module then evaluates each health condition, following a structured, step-wise approach to determine eligibility and make the final matching decision. This paper enhances patient-trial matching by leveraging the agency and reasoning capabilities of large language models (LLMs) through automated agent interactions, including collaboration, critique, and navigation. Experimental results on a public dataset demonstrate that our framework surpasses existing benchmarks, achieving up to an 8% improvement in accuracy for specific criteria. Furthermore, in a real-world offline test, MAKAR achieved a 100% accuracy. These findings show MAKAR’s potential as a scalable and robust solution for clinical trial patient matching.
Patient Matching, Multi-agent Systems, Knowledge Augmentation, Reasoning LLMs
ccs: Computing methodologies Multi-agent systemsccs: Applied computing Health care information systems
1. Introduction
Patient matching for clinical trials refers to the process of identifying and enrolling participants whose health profiles meet the specific eligibility criteria of a given clinical study. Efficient patient matching is crucial to accelerating the drug development process and reducing the overall cost of clinical trials. It also provides an opportunity for participants to access experimental treatments that could potentially improve their health outcomes and, in some cases, transform their quality of life.
@@ 1. Introduction
The emergence of of LLMs offers a promising opportunity to improve patient matching. LLMs, such as GPT-4 (Achiam et al., 2023), usually possess a large text window and advanced natural language understanding capabilities (Karanikolas et al., 2023), which enable them process complex trial criteria and patient data effectively. However, their general knowledge may not always be sufficient for knowledge-intensive tasks like patient matching in clinical trials, where domain-specific expertise is required. The knowledge gap inherent in LLMs and misssing information in the pre-defined criteria can limit their performance, highlighting the need for LLM augmentation of specialized medical knowledge (Shuster et al., 2021). While research indicates that LLMs do encode medical knowledge in a parameterized form (Singhal et al., 2023), this information can be outdated or misaligned with the special needs of certain clinical trials.
Figure 1. The workflow of MAKAR framework
@@
To address these limitations, we present a new framework called: Multi-Agents for Knowledge Augmentation and Reasoning (MAKAR). MAKAR is a novel framework designed to enhance the patient matching process. The framework consists of two core modules: the Augmentation Module and the Reasoning Module. The Augmentation Module enhances trial criteria by providing detailed explanations and includes three key components: a Router Agent, multiple Augmentation Agents, and a Critic Agent. The Reasoning Module systematically evaluates each criterion-related condition in a step-by-step manner to determine eligibility and make the final matching decision. It comprises a Reasoning Agent and a Supervision Matching Agent. By leveraging this multi-agent framework, MAKAR improves the performance on patient-trial matching by dynamically incorporating domain-specific knowledge and step-wise reasoning. Our experimental results demonstrate the effectiveness of MAKAR. On a public dataset, our framework achieves an average improvement of 3% over benchmark methods, with up to an 8% improvement for complex criteria. Furthermore, in an offline real-world data test, MAKAR outperformed the state-of-the-art closed-source reasoning LLMs and achieved 100% accuracy, showing its reliability in practical applications. Our key contributions are as follows:
  • We propose an automated method to refine and enhance trial criteria, making them more interpretable and suitable for AI-driven patient-trial matching.
  • We introduce a multi-agent framework that combines augmentation and reasoning, achieving strong performance across both public and real-world datasets.
  • We identify a gap in existing approaches and highlight the importance of explicitly incorporating well-structured, detailed explanations of related concepts to improve LLM-driven clinical reasoning.
2. Related Work
2.1. Clinical Trial Patient Matching
Traditional patient matching systems rely on query-based structured databases, such as Criteria2Query (Yuan et al., 2019), or handcrafted rule-based approaches like RBC (Oleynik et al., 2019). However, these methods struggle to handle unstructured free-text data and are highly sensitive to variations in formatting and verbal expressions. Unfortunately, a significant portion of EHRs consists of free-text, including clinical notes, surgical records, radiology reports, pathology reports and so on (Sun et al., 2018). Thus, the effectiveness and scalability of traditional methods are significantly limited, making them unsuitable for future patient matching tasks.
Recent advancements in LLMs have shown remarkable capabilities in directly processing free-text clinical trial criteria, offering scalable solutions without the need for extensive rule-based preprocessing. Many Studies have applied these models to patient-centric trial matching tasks using real-world data. For instance, Nievas et al. (2024) fine-tuned the open-source Llama 2 (Touvron et al., 2023) to automate patient matching. However, fine-tuning LLMs is computationally expensive, prompting alternative approaches that leverage frozen models. Jin et al. (2023) introduced TrialGPT, a patient-centric matching framework that assesses patient eligibility on a criterion-by-criterion basis, filtering out ineligible trials and ranking potentially eligible ones. Their results indicated a strong correlation between TrialGPT’s predictions and expert clinician judgments, achieving accuracy at the criterion level similar to human experts. Similarly, Gupta et al. (2024) proposed OncoLLM, an end-to-end pipeline that converts trial criteria into conversational prompts and uses a Chain-of-Thought (CoT) (Wei et al., 2022) reasoning strategy. OncoLLM outperforms GPT-3.5 while underperforms GPT-4, despite its smaller model size.
In contrast, other studies have focused on trial-centric patient matching, typically using public deidentified datasets due to HIPAA-related data constraints. Wornow et al. (2024) developed a LLM-based system with a two-stage retrieval pipeline to identify suitable patients for trials using a zero-shot prompt approach. It reduces token processing costs while outperforming existing state-of-the-art models on the n2c2 2018 cohort selection benchmark (Stubbs et al., 2019). Likewise, Beattie et al. (2024) applied a CoT strategy in a Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) system using GPT-3.5 Turbo and GPT-4, demonstrating results comparable to those of Wornow et al. (2024).
These works collectively highlight the potential of LLMs in clinical trial matching. However, existing approaches directly employ LLMs without verifying whether the trial criteria are sufficiently understood by the model (Webson and Pavlick, 2021) or whether there are gaps in the model’s knowledge (Lewis et al., 2020). Our work fits within the trial-centric patient matching literature and makes a novel contribution by automatically identifying and addressing these gaps, thereby enhancing the accuracy patient matching.
2.2. LLM-based Multi-Agents (LLM-MA)
LLM-MA have seen rapid development across various domains, driven by the promising potential of LLMs as autonomous agents.  Guo et al. (2024) provide a comprehensive overview of recent advancements in LLM-MA applications, categorizing them into two primary areas: Problem-Solving and World Simulation. Problem-solving LLM-MA systems explore the capabilities of agents with specialized expertise, enabling them to outperform single-agent LLM systems through collaborative efforts. For instance,  Wang et al. (2024) proposed Mixture-of-Agents, a layered multi-agent architecture that achieves state-of-the-art performance in benchmarks like AlpacaEval 2.0 (Dubois et al., 2024).
In the healthcare domain, the application of LLM-MA systems is rapidly gaining traction. Yue et al.(Yue and Fu, 2024) introduce ClinicalAgent, a multi-agent architecture powered by LLM reasoning for multi-agent collaboration, achieving higher precision in predicting clinical trial outcomes. Similarly, Liao et al.(Liao et al., 2024) propose REFLECTool, a multi-agent framework that enables agents to learn how to select domain-specific tools, thereby extend their capabilities. TriageAgent(Lu et al., 2024) leverages LLMs for role-playing interactions, integrating self-confidence assessments and early-stopping mechanisms in multi-round discussions to improve document reasoning and classification accuracy for triage tasks. Tang et al.(Tang et al., 2023) use multiple LLM-based agents as domain experts to collaboratively analyze medical reports. And Li et al. (Li et al., 2024) further introduce the multi-agent collaboration framework to handle multimodal tasks.
To the best of our knowledge, this study is the first to apply an LLM-MA framework for cohort selection in clinical trials. Our approach explore the agency and reasoning capabilities of LLM-MA to dynamically enhance patient-trial matching with tailored knowledge. Our approach allows a better contextual understanding of patient profiles and lead to improved matching performance and more informed decision-making in clinical trial recruitment.
Criteria Accuracy Precision Recall F1-score
Zero-shot CoT MAKAR Zero-shot CoT MAKAR Zero-shot CoT MAKAR Zero-shot CoT MAKAR
ABDOMINAL 0.858 0.868 0.847 0.893 0.897 0.944 0.701 0.729 0.626 0.785 0.804 0.753
ADVANCED-CAD 0.854 0.854 0.868 0.895 0.872 0.929 0.853 0.882 0.841 0.873 0.877 0.883
ALCOHOL-ABUSE 0.969 0.969 0.976 0.529 0.545 0.667 0.900 0.600 0.600 0.667 0.571 0.632
ASP-FOR-MI 0.861 0.875 0.885 0.899 0.908 0.919 0.930 0.939 0.939 0.915 0.923 0.929
CREATININE 0.802 0.802 0.885 0.664 0.667 0.812 0.934 0.925 0.896 0.776 0.775 0.852
DIETSUPP-2MOS 0.764 0.781 0.785 0.704 0.717 0.721 0.940 0.953 0.953 0.805 0.818 0.821
DRUG-ABUSE 0.941 0.920 0.993 0.469 0.395 0.882 1.000 1.000 1.000 0.638 0.566 0.938
ENGLISH 0.976 0.986 0.983 0.985 0.989 0.996 0.989 0.996 0.985 0.987 0.992 0.991
HBA1C 0.840 0.830 0.878 0.722 0.702 0.807 0.892 0.902 0.863 0.798 0.790 0.834
KETO-1YR 0.993 0.990 0.993 - - - - - - - - -
MAJOR-DIABETES 0.833 0.826 0.847 0.851 0.831 0.873 0.840 0.853 0.840 0.845 0.842 0.856
MAKES-DECISIONS 0.920 0.924 0.931 0.985 0.985 0.996 0.931 0.935 0.931 0.957 0.959 0.963
MI-6MOS 0.885 0.872 0.927 0.434 0.396 0.571 0.885 0.808 0.769 0.582 0.532 0.656
Average 0.884 0.884 0.908 0.750 0.740 0.841 0.899 0.876 0.853 0.801 0.786 0.841
Table 1. Criterion-Level Comparison Across Different Methods for N2C2
3. Methods
@@
The proposed framework, shown in Figure 1, consists of two core modules: the Augmentation Module and the Reasoning Module. The augmentation module is designed to generate criteria with comprehensive explanations of all relevant concepts, while the reasoning module evaluates each condition within the criteria to determine eligibility and make the final matching decision. Additionally, the framework contains five interconnected components, each outlined below.
@@
Router Agent In the MAKAR framework, multiple specialized augmentation agents are available to augment the criterion with domain-specific knowledge. The Router Agent determines the most suitable augmentation agent for the task and routes the workflow accordingly. Inspired by SELFREF (Kadavath et al., 2022) and MOREINFO (Feng et al., 2023), we adopt a prompt-based self-probing strategy. If the agent determines that the original criterion is detailed and understandable enough for other agents, the criterion will be directly forwarded to the Reasoning Module. Otherwise, the criterion is forwarded to the Knowledge Augmentation Agent chosen by the Router Agent for further processing.
@@
Knowledge Augmentation Agent: This component augments the criterion with relevant, domain-specific information from different source. For instance, the Retrieval Agent (Lewis et al., 2020) retrieves information from indexed medical databases, the Online Search Agent gathers supplementary knowledge from online sources, and the Self-Augment Agent leverages the LLM’s intrinsic knowledge to refine the original criterion.
@@
Critic Agent: Acting as a supervisor, this agent ensures that the augmented criterion is aligned with the original instructions. For instance, the original criterion must be preserved verbatim. Additionally, the critic agent verifies that the augmented output is well-structured (e.g., concepts are clearly itemized) to facilitate further step-wise reasoning. This critic mechanism are designed to improve the framework’s ability to follow instructions accurately while maintaining consistency and clarity.
@@
Reasoning Agent With the augmented criterion and patient EHR data, the Reasoning Agent evaluates patient eligibility by assessing for each concepts mentioned in the augmented criterion. It generates a detailed reasoning summary, including assessments for all relevant health conditions, and forwards it to the Supervision Matching Agent.
@@
4. Evaluation & Results
Due to the sensitive nature of patient data, which is strictly protected under The Health Insurance Portability and Accountability Act (HIPAA), and the compliance requirements for processing detail clinical trial documents, full-scale online deployment is not currently feasible—particularly due to concerns about data security when using closed-source LLMs. Instead, to evaluate the effectiveness of our methods, we conducted two experiments: (1) an assessment using a public dataset and (2) a limited-scale offline evaluation with deidentified real-world data.
4.1. Data
@@
The public dataset we use is from Track 1 of the 2018 n2c2 challenge (Stubbs et al., 2019), which is commonly used for evaluating patient matching in clinical trials. The dataset consists of 288 de-identified patients, each having between 2 to 5 unstructured longitudinal clinical notes in American English. The entire corpus contains 781,006 tokens, with an average of 2,711 tokens per patient. The challenge simulates a synthetic clinical trial with 13 predefined inclusion criteria, resulting in 3744 patient-criteria pairs in total. The detailed criteria can be found in Table 6. For each patient-criteria pair, a binary label is provided, indicating whether the patient ”MET” or ”NOT MET” the criterion. Two annotators with medical background labeled these data. The disagreements between the two annotators were resolved by an independent MD to establish the final ground truth labels.
@@
The criteria in the n2c2 dataset are relatively straightforward and the patient records are well-structured, making them easy to follow. To evaluate the potential of our proposed method in a more complex scenario, we constructed a real-world dataset. Specifically, we sampled 30 eligibility criteria—15 inclusion and 15 exclusion—from ongoing breast cancer trials listed on ClinicalTrials.gov. We then manually selected 10 patients who have been diagnosed with breast cancer from our in-house EHR system. For each patient, we extracted their most recent deidentified clinical visit record. An expert subsequently asseessed whether each patient met each criterion, resulting in 300 labeled criterion-patient pairs. Notably, if a patient’s information was insufficient to determine eligibility for a given criterion, they were labeled as ”NOT MET” to ensure a conservative patient selection. We listed the sampled criterion in Table  9
4.2. Baselines and Metrics
Previous research has demonstrated the superiority of LLM-based methods over traditional rule-based systems; hence, this study focuses exclusively on LLM-based benchmarks. We include two approaches in Experiment 1: the zero-shot method with hand-crafted criteria (Wornow et al., 2024), as shown in Table LABEL:table:redefined_criteria, and a Chain-of-Thought (CoT) approach (Beattie et al., 2024) using the original criteria. In addtion, we also include o3-mini from OpenAI as the benchmark in Experiment 2 since reasoning LLMs have shown significant improvement on complex clinical scenarios(Xie et al., 2024). In line with prior literature, we report accuracy, precision, recall, and F1-score for this binary classification task.
Previous studies using n2c2 dataset have primarily focused on evaluating classification performance at the individual criterion level, often overlooking the overall patient-trial assessment. In this study, we also examine patient-trial selection performance in Experiment 1. Specifically, a patient is eligible for the trial only if all inclusion criteria are met. However, no patient in the n2c2 dataset satisfied all 13 criteria. To mitigate this, we synthesize a trial with selected criteria that had at least 100 patients meeting them, resulting in 28 eligible patients from a total cohort of 288. We then evaluated our method’s on this new synthetic trial in identifying eligible patients and reported the findings.
4.3. Experiment Settings

### v3

@@
Abstract
Matching patients effectively and efficiently for clinical trials is a significant challenge due to the complexity and variability of patient profiles and trial criteria. This paper introduces Multi-Agent for Knowledge Augmentation and Reasoning (MAKAR), a novel multi-agent system that enhances patient-trial matching by integrating criterion augmentation with structured reasoning. MAKAR consistently improves performance by an average of 7% across different datasets. Furthermore, it enables privacy-preserving deployment and maintains competitive performance when using smaller open-source models. Overall, MAKAR can contributes to more transparent, accurate, and privacy-conscious AI-driven patient matching.
Enhancing Clinical Trial Patient Matching through Knowledge Augmentation and Reasoning with Multi-Agent
Hanwen Shi University of Maryland hwshi@umd.edu                        Jin Zhang Pi Health Jin.Zhang@pihealth.ai                        Kunpeng Zhang University of Maryland kpzhang@umd.edu
Figure 1: Comparison of accuracy of zero-shot patient matching with different prompts. The values are averaged over 5 runs; error bars show the observed range.
1 Introduction
Patient matching for clinical trials is the process of identifying and enrolling participants whose health profiles meet the specific eligibility criteria of a given clinical study. Efficient patient matching is crucial to accelerating the drug development process and reducing the overall cost of clinical trials. It also provides an opportunity for participants to access experimental treatments that could potentially improve their health outcomes and, in some cases, transform their quality of life.
@@ 1 Introduction
The emergence of of LLMs offers a promising opportunity to improve patient matching. LLMs, such as GPT-4 Achiam et al. (2023), usually possess a large text window and advanced natural language understanding capabilities (Karanikolas et al., 2023), which enable them process complex trial criteria and patient data effectively. However, three limitations constrain the performance of LLMs. First, knowledge gaps (Shuster et al., 2021) may exist due to training data cut-off dates and incomplete coverage of specialized medical knowledge, such as newly approved drugs, treatments, or updated FDA guidelines. Second, hallucinations (Huang et al., 2025) remain a significant challenge: LLMs are often prompted to generate detailed reasoning to improve interpretability and trustworthiness in patient matching (Jin et al., 2024), yet such reasoning is unverified and potentially misleading. Finally, prompt sensitivity can substantially impact outputs, as even minor variations in prompts can lead to divergent results (Zhuo et al., 2024).
@@
To address these limitations, we present a training-free multi-agent workflow: Multi-Agent for Knowledge Augmentation and Reasoning (MAKAR). MAKAR consists of two multi-agent modules: the Augmentation Module and the Reasoning Module. The augmentation module enriches trial criteria by providing detailed explanations from different sources. And the reasoning module evaluates each criterion-related patient condition in a step-by-step manner to determine eligibility and make the final matching decision. We demonstrate the effectiveness of MAKAR our framework on public dataset and in-house real-world database. On average, MAKAR achieves an improvement of 7% over benchmark methods, with up to a 10% improvement on certain criteria. Our key contributions are as follows:
  • We demonstrate by experiment that LLMs are sensitive to how criteria are formulated.
  • We introduce MAKAR, a double-module multi-agent workflow that allows complementary contributions from criterion augmentation and reasoning. And it achieves consistent performance improvement across different datasets.
  • We demonstrate that MAKAR remains effective when deployed with open-source language models. And it has potential to contribute to more trustworthy and privacy-preserving AI application in drug development.
2 Criterion Sensitivity
To further motivate this work and demonstrate existing gaps, we conducted a preliminary experiment comparing patient matching performance on the 2018 n2c2 Challenge Track 1 dataset (Stubbs et al., 2019) using three different criteria: (1) the original criteria provided by the challenge (see Appendix B.1), (2) human-refined criteria used by Wornow et al. (2024) (Appendix B.2), and (3) GPT-4o-refined criteria (Appendix B.3). All experiments were performed with GPT-4o in a zero-shot setting, and the prompts are provided in Appendix C.1. The prompt used to refine the criteria in (3) is provided in Appendix C.2. The results are presented in Figure 1.
Overall, the results in Figure 1 suggests three important observations. First, the criterion sensitivity of LLM-based patient matching is substantial: simply varying the formulation of the criteria can cause notable differences in accuracy across most conditions. Second, while using human-refined criteria does not universally outperform the original criteria, automatically refined and augmented prompts generated by GPT-4o consistently improve performance. Third, explicitly incorporating detailed knowledge (e.g., explanations) into the prompt can further enhance accuracy, even when the LLMs themselves do not exhibit an obvious knowledge gap. For example, for the "CREATININE" criterion, GPT-4o was able to return the correct reference ranges by gender and age when being asked directly. However, explicitly adding these ranges in the prompt—as done in the GPT-4o-augmented criteria—increased performance by 8% compared to prompts that used only general descriptions such as “normal range.” These findings highlight the need for methods that can systematically augment knowledge rather than relying solely on prompt engineering or implicit model knowledge.
Figure 2: The workflow of MAKAR framework
3 Methods
@@
Our proposed methods, shown in Figure 2, is a training-free multi-agent workflow and consists of two core modules: the Augmentation Module and the Reasoning Module. The augmentation module is designed to generate criteria with comprehensive explanations of relevant concepts, while the reasoning module evaluates each condition within the criteria to determine eligibility and make the final matching decision. The workflow contains five interconnected components and all prompts used by the agents listed below can be found in Appendix C.3. The pseudo code of the workflow is provided in Appendix D
@@
Router Agent In MAKAR, multiple specialized augmentation agents are available to augment the criterion with domain-specific knowledge. The Router Agent determines the most suitable augmentation agent for the task and routes the workflow accordingly. Inspired by SELFREF Kadavath et al. (2022) and MOREINFO Feng et al. (2023), we adopt a prompt-based self-probing strategy. If the agent determines that the original criterion is detailed and understandable enough for other agents, the criterion will be directly forwarded to the Reasoning Module. Otherwise, the criterion is forwarded to the Knowledge Augmentation Agent chosen by the Router Agent for further processing.
@@
Knowledge Augmentation Agent: This component augments the criterion with relevant, domain-specific information from different source. For instance, the Retrieval Agent Lewis et al. (2020) retrieves information from indexed medical databases, the Online Search Agent gathers supplementary knowledge from online sources, and the Self-Augment Agent leverages the LLM’s intrinsic knowledge to refine the original criterion.
@@
Critic Agent: We add Critic Agent as the supervisor to power the iterative refinement with self-feedback as in SELF-REFINE Madaan et al. (2023). It ensures that the augmented criterion is aligned with the original instructions. For instance, the original criterion must be preserved verbatim. Additionally, the critic agent verifies that the augmented output is well-structured (e.g., concepts are clearly itemized) to facilitate further step-wise reasoning. This critic mechanism are designed to improve the framework’s ability to follow instructions accurately while maintaining consistency and clarity.
@@
Reasoning Agent With the augmented criterion and patient EHR data, the Reasoning Agent evaluates patient eligibility by assessing for each concepts mentioned in the augmented criterion. It generates a detailed reasoning summary Jin et al. (2024), including assessments for all relevant health conditions, and forwards it to the Matching Agent.
@@
Criteria Jin et al. (2024) Wornow et al. (2024)  Beattie et al. (2024) MAKAR
Accuracy F1 Accuracy F1 Accuracy F1 Accuracy F1
N2C2 Dataset
ABDOMINAL $0.813$ $0.722$ $0.858$ $0.785$ 0.868 0.804 $0.809$ $0.671$
ADVANCED-CAD $0.823$ $0.859$ $0.854$ $0.873$ $0.854$ $0.877$ 0.882 0.897
ALCOHOL-ABUSE $0.972$ $0.667$ $0.969$ $0.667$ $0.969$ $0.571$ 0.983 0.706
ASP-FOR-MI 0.906 0.944 $0.861$ $0.915$ $0.875$ $0.923$ $0.896$ $0.938$
CREATININE $0.802$ $0.776$ $0.802$ $0.776$ $0.802$ $0.776$ 0.924 0.900
DIETSUPP-2MOS $0.722$ $0.759$ $0.764$ $0.805$ $0.781$ $0.818$ 0.826 0.846
DRUG-ABUSE $0.882$ $0.469$ $0.941$ $0.638$ $0.920$ $0.566$ 0.972 0.765
ENGLISH $0.948$ $0.971$ $0.976$ $0.987$ $0.986$ $0.992$ 0.997 0.998
HBA1C $0.840$ $0.800$ $0.840$ $0.798$ $0.830$ $0.790$ 0.934 0.899
KETO-1YR 0.993 - 0.993 - $0.990$ 0.993 -
MAJOR-DIABETES $0.830$ $0.838$ $0.833$ 0.845 $0.826$ $0.842$ 0.840 $0.838$
MAKES-DECISIONS $0.580$ $0.722$ $0.920$ $0.957$ $0.924$ $0.959$ 0.965 0.982
MI-6MOS $0.833$ $0.510$ $0.885$ $0.582$ $0.872$ $0.532$ 0.962 0.807
Average $0.842$ $0.753$ $0.884$ $0.802$ $0.884$ $0.787$ 0.922 0.854
ClinicalTrial Dataset
CT-IN $0.900$ $0.828$ $0.967$ $0.929$ 0.987 0.982
CT-EX $0.900$ $0.885$ $0.947$ $0.937$ 0.987 0.982
Table 1: Performance Comparison Across Models on N2C2 and ClinicalTrial Datasets. CT-IN and CT-EX stand for inclusion and exclusion criteria in the ClinicalTrial dataset. The F1 score for KETO-1YR is missing because only one patient met this criterion, and no method match it correctly.
4 Experiments
4.1 Data
@@
We use a widely used public dataset and an in-house real-world dataset to evaluate our proposed method. Both datasets were labeled by at least two trained annotators, with any disagreements resolved by an independent physician to determine the final ground truth. The public dataset is from Track 1 of the 2018 n2c2 challenge Stubbs et al. (2019). The dataset consists of 288 de-identified patients, each having between 2 to 5 unstructured longitudinal clinical notes in American English. The challenge simulates a synthetic clinical trial with 13 predefined inclusion criteria (see Appendix B.1), resulting in 3744 patient-criteria pairs. Each pair has a binary label—“MET” or “NOT MET” and the statistics can be found in Table 4.
@@
To assess our method in a more complex setting, we created a real-world dataset by sampling 30 eligibility criteria (15 inclusion, 15 exclusion) from ongoing breast cancer trials on ClinicalTrials.gov. We manually selected 10 breast cancer patients from our in-house EHR system and extracted their most recent deidentified clinical records. When information was insufficient to confirm eligibility, the criterion was conservatively labeled as "NOT MET." The sampled criteria are listed in Appendix B.4.
4.2 Baselines and Metrics
Previous research has shown that LLM-based approaches outperform traditional rule-based systems Jin et al. (2024). Thus, this study focuses exclusively on LLM-based benchmarks. We include three benchmarks: (1) TrialGPT Jin et al. (2024), (2) the two-stage retrieval-based zero-shot matching system from Wornow et al. (2024), and (3) the chain-of-thought (CoT) based method introduced by Beattie et al. (2024). For the n2c2 dataset, we report accuracy and F1-score for each criterion as well as the overall average. For the ClinicalTrial dataset, we report the average performance for inclusion and exclusion criteria separately.
4.3 Experiment Settings